# Online Convex Optimization with Heavy Tails: Old Algorithms, New Regrets, and Applications

### v1

@@ Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
Equipped with Theorem 10, we show the following anytime last-iterate convergence rate for $\mathsf{SGD}$$\mathsf{OGD}$. As far as we know, this is the first and the only provable result demonstrating that the last iterate of $\mathsf{SGD}$ can converge in heavy-tailed stochastic optimization without gradient clipping. Compared to Corollary 2, the difference is up to an extra logarithmic factor. Therefore, it is nearly optimal.
Corollary 5.
Under Assumption 1 for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$, taking $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ in $\mathsf{OGD}$ (Algorithm 1), we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}},\forall\boldsymbol{x}\in\mathcal{X}.$
Proof.
By Theorem 10, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]$
$\lesssim$ $\frac{D^{2}}{\sum_{t=1}^{T}\eta_{t}}+G^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=(t+1)\land T}^{T}\eta_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T}\frac{\eta_{t}^{\mathsf{p}}}{\sum_{s=(t+1)\land T}^{T}\eta_{s}}$
$=$ $\frac{D^{2}}{\sum_{t=1}^{T}\eta_{t}}+G^{2}\left(\eta_{T}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}}{\sum_{s=t+1}^{T}\eta_{s}}\right)+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(\eta_{T}^{\mathsf{p}-1}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}}{\sum_{s=t+1}^{T}\eta_{s}}\right).$
For any $t\in\left\{0\right\}\cup\left[T-1\right]$, observe that by Cauchy-Schwarz inequality
$(T-t)^{2}\leq\left(\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}\right)\left(\sum_{s=t+1}^{T}\eta_{s}\right)\Rightarrow\frac{1}{\sum_{s=t+1}^{T}\eta_{s}}\leq\frac{\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}}{(T-t)^{2}}.$
Thus, there is
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim$ $\frac{D^{2}}{T^{2}}\sum_{t=1}^{T}\frac{1}{\eta_{t}}+G^{2}\left(\eta_{T}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}}{(T-t)^{2}}\right)$
$+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(\eta_{T}^{\mathsf{p}-1}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}}{(T-t)^{2}}\right).$ (33)
We first bound
$\sum_{t=1}^{T}\frac{1}{\eta_{t}}=\sum_{t=1}^{T}\frac{G\sqrt{t}}{D}\lor\frac{\sigma t^{1/\mathsf{p}}}{D}\leq\sum_{t=1}^{T}\frac{G\sqrt{t}}{D}+\frac{\sigma t^{1/\mathsf{p}}}{D}\lesssim\frac{G}{D}T^{3/2}+\frac{\sigma}{D}T^{1+1/\mathsf{p}},$
@@
$\frac{D^{2}}{T^{2}}\sum_{t=1}^{T}\frac{1}{\eta_{t}}\lesssim\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}.$ (34)
Next, we know
$\eta_{T}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}}{(T-t)^{2}}$ $\overset{(a)}{\leq}\frac{D}{G\sqrt{T}}+\sum_{t=1}^{T-1}\left[\frac{D}{G}\cdot\frac{\sum_{s=t+1}^{T}\sqrt{s}}{t(T-t)^{2}}+\frac{\sigma D}{G^{2}}\cdot\frac{\sum_{s=t+1}^{T}s^{1/\mathsf{p}}}{t(T-t)^{2}}\right]$
@@
where $(a)$ is by $\eta_{t}\leq\frac{D}{G\sqrt{t}}$ and $\frac{1}{\eta_{s}}\leq\frac{G\sqrt{s}}{D}\lor\frac{\sigma s^{1/\mathsf{p}}}{D}$. Hence, there is
$G^{2}\left(\eta_{T}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{2}\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}}{(T-t)^{2}}\right)\lesssim\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}}.$ (35)
Similarly, we can bound
$\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(\eta_{T}^{\mathsf{p}-1}+\sum_{t=1}^{T-1}\frac{\eta_{t}^{\mathsf{p}}\sum_{s=t+1}^{T}\frac{1}{\eta_{s}}}{(T-t)^{2}}\right)\lesssim\frac{GD\left(1+\log T\right)}{\sqrt{T}}+\frac{\sigma D\left(1+\log T\right)}{T^{1-\frac{1}{\mathsf{p}}}}.$ (36)
Finally, we plug (34), (35) and (36) back into (33) to conclude. ∎
@@ E.6.2 Useful Existing Results
Note that Lemma 4 can be only applied to a randomzied algorithm with bounded outputs. To overcome this issue, we need the following variants of $\bar{F}_{T,U}$ and $\bar{\boldsymbol{g}}_{T,U}$ introduced by [7]:
@@
    Let $\boldsymbol{u}_{j_{t}}\in\mathbb{R}^{d}$ denotes the $j_{t}$-th column of $U$. By the definiton of $j_{t}$, there is
@@ E.6.3 Analysis under Heavy-Tailed Noise
The instances $\widehat{F}_{T,U}$ and $\widehat{\boldsymbol{g}}_{T,U}$ (see (49) and (50)) constucted based on $F_{T}$ in Lemma 7 and $\boldsymbol{g}_{T}$ in Lemma 8 under $\eta=\frac{1}{\sqrt{5}}-\frac{39}{140}$ and $R=7\gamma\sqrt{T}$ for $\gamma=23$ satisfy the following properties:
@@ F.1 Full Results for Smooth $\ell_{t}$
When $\boldsymbol{x}^{\star}\in\mathcal{X}$ is also a global minimizer (i.e., $\boldsymbol{x}^{\star}\in\mathrm{arginf}_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})$), we have $\left\|\nabla F(\boldsymbol{x}^{\star})\right\|=0$ since $F$ is differentiable. This fact recovers Theorem 7.
@@ F.2.2 New Regret for $\mathsf{OAdaGrad}$
Equipped with Lemma 10, we first prove the following Theorem 14, which establishs the regret of $\mathsf{OAdaGrad}$ under heavy tails.
@@
To gain some intuition of the new Condition 2, we first let $H_{t}=0$ to have $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim G_{t}$, which is similar to $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{y})\right\|\lesssim G_{t},\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$ in Assumption 3. Next, we consider the case $G_{t}=0$, where Condition 2 states an inequality $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim H_{t}^{\frac{1}{1+\nu}}(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})^{\frac{\nu}{1+\nu}}$, which is known to hold if $\ell_{t}$ is $(H_{t},\nu)$-Hölder smooth on $\mathbb{R}^{d}$ and convex, which can be viewed as a generalization of Assumption 3 (though strictly speaking, we only require $\ell_{t}$ to be Hölder smooth on $\mathcal{X}$ in Assumption 3). For general $G_{t}$, $H_{t}$, and $\nu$, whether Condition 2 is strictly general than the thrird point in Assumption 3 is unclear. Thus, we consider it as a separate condition here.
@@ G.2.2 Convergence Rates
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\left\|\nabla F(\boldsymbol{y}_{0})\right\|}{KT}+\frac{F(\boldsymbol{y}_{0})-F_{\star}}{\delta K}+\frac{H\delta^{\nu}}{T^{\frac{1}{2}+\nu}}+\frac{G}{\sqrt{T}}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}.$
Similar to before, we can consider two situations where problem-dependent parameters are known or unknown, corresponding to the following Corollaries 9 and 10, respectively. To save space, the proofs are omitted since they can be easily checked.
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{y}}_{k})\right\|_{\delta}\right]\lesssim$ $\frac{\nabla}{N}+\frac{H\delta^{\nu}}{N^{\frac{1}{2}+\nu}}+\frac{G}{\sqrt{N}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\Delta}{\delta N}$
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{y}}_{k})\right\|_{\delta}\right]\lesssim$ $\frac{\nabla}{N}+\frac{H\delta^{\nu}}{N^{\frac{1}{2}+\nu}}+\frac{H}{\delta^{\frac{2-\nu}{5}}N^{\frac{1+2\nu}{5}}}$
@@ G.2.3 A Special Case: Hölder Smooth Nonconvex Functions
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|\right]\lesssim\frac{\nabla}{KT}+\frac{\Delta}{\delta K}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}+H\delta^{\nu}.$
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{y}}_{k})\right\|\right]\lesssim\frac{\nabla}{N}+\frac{H^{\frac{1}{1+\nu}}\Delta^{\frac{\nu}{1+\nu}}}{N^{\frac{\nu}{1+\nu}}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\sigma^{\frac{\mathsf{p}\nu}{r}}H^{\frac{\mathsf{p}-1}{r}}\Delta^{\frac{(\mathsf{p}-1)\nu}{r}}}{N^{\frac{(\mathsf{p}-1)\nu}{r}}}.$
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{y}}_{k})\right\|\right]\lesssim\frac{\nabla}{N}+\frac{\Delta}{N^{\frac{1}{4}}}+\frac{H}{N^{\frac{\nu}{4}}}+\frac{\sigma}{N^{\frac{\mathsf{p}-1}{2\mathsf{p}}}}.$
@@
For general $\nu\in\left(0,1\right)$, as far as we know, no previous works consider heavy-tailed noise. Hence, both these Corollaries are the first and new.

### v2

@@ Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
$=\frac{2HD^{2}}{T-1}+\frac{G^{2}}{\mu}\sum_{t=1}^{T-1}\frac{1}{(T-t)(t+2\kappa)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\sum_{t=1}^{T-1}\frac{1}{(T-t)(t+2\kappa)^{\mathsf{p}-1}}.$ (41)
Given $a\in\left(0,1\right]$, we can bound
$\sum_{t=1}^{T-1}\frac{1}{(T-t)(t+2\kappa)^{a}}=\frac{1}{T+2\kappa}\sum_{t=1}^{T-1}\frac{(t+2\kappa)^{1-a}}{T-t}+\frac{1}{(t+2\kappa)^{a}}\lesssim\frac{1+\log T}{(T+2\kappa)^{a}},$
where the last step is due to
$\sum_{t=1}^{T-1}\frac{(t+2\kappa)^{1-a}}{T-t}\lesssim(T+2\kappa)^{1-a}(1+\log T)$ and $\sum_{t=1}^{T-1}\frac{1}{(t+2\kappa)^{a}}\lesssim\begin{cases}T^{1-a}&a\in\left(0,1\right)\\ 1+\log T&a=1\end{cases}.$
Finally, we apply the above inequality to (41) with $a=1$ and $a=\mathsf{p}-1$ to obtain
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{T}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}}.$
For $\eta_{t}=\begin{cases}\frac{1}{\mu(1+2\kappa)}&t\leq\tau\\ \frac{2}{\mu(t-\tau+2+4\kappa)}&t\geq\tau+1\end{cases}$, let $\eta\triangleq\frac{1}{\mu(1+2\kappa)}$, we can find
$\gamma_{t}=\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}=\begin{cases}\frac{\eta}{(1-\mu\eta)^{t-1}}&t\leq\tau\\ \frac{\eta(t-\tau+1+4\kappa)}{(1-\mu\eta)^{\tau-1}(1+4\kappa)}&t\geq\tau+1\end{cases}.$ (42)
Hence, there is
$\frac{1-\mu\eta_{1}}{\sum_{t=1}^{T-1}\gamma_{t}}$ $\leq\frac{1-\mu\eta_{1}}{\sum_{t=1}^{\tau}\gamma_{t}}=\frac{1-\mu\eta}{\sum_{t=1}^{\tau}\frac{\eta}{(1-\mu\eta)^{t-1}}}=\frac{(1-\mu\eta)^{\tau}}{\eta\sum_{t=1}^{\tau}(1-\mu\eta)^{\tau-t}}=\frac{\mu(1-\mu\eta)^{\tau}}{1-(1-\mu\eta)^{\tau}}$
$\overset{(1-\mu\eta)^{\tau}\leq 1-\mu\eta}{\leq}\frac{(1-\mu\eta)^{\tau}}{\eta}=2H(1-\mu\eta)^{\tau-1}\leq 2H\exp\left(-\mu\eta(\tau-1)\right)$
$=2H\exp\left(\frac{-\tau+1}{1+2\kappa}\right)\overset{\kappa\geq 0}{\leq}2eH\exp\left(\frac{-\tau}{1+2\kappa}\right)\overset{\tau\geq\frac{T}{2}}{\lesssim}\frac{H}{\exp\left(\frac{T}{2+4\kappa}\right)},$
implying that
[…]
Given $a\in\left(0,1\right]$, we can write
$\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}=\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sum_{t=\tau+1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}=\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sum_{t=\tau+1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}.$ (44)
When $t\leq\tau$, we know
[…]
@@
$\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}\leq\frac{2(1+4\kappa)\tau}{\mu^{a}(1+2\kappa)^{a}(T-\tau)(T-\tau+1+8\kappa)}\lesssim\frac{(1+\kappa)^{1-a}}{\mu^{a}(T+\kappa)}\leq\frac{1}{\mu^{a}(T+\kappa)^{a}}.$ (45)
When $t\geq\tau+1$, we know
@@
$=\frac{2}{2T-2\tau+1+8\kappa}\left(\frac{t-\tau+1+4\kappa}{T-t}+\frac{t-\tau+1+4\kappa}{T+t-2\tau+1+8\kappa}\right)$
$\overset{\tau\leq\frac{T+1}{2},t\geq\tau+1}{\lesssim}\frac{1}{T+\kappa}\left(\frac{t-\tau+1+4\kappa}{T-t}+\frac{t-\tau+1+4\kappa}{T-\tau+2+8\kappa}\right),$
which implies
$\sum_{t=\tau+1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}$ $\lesssim\frac{1}{\mu^{a}(T+\kappa)}\sum_{t=\tau+1}^{T-1}\frac{(t-\tau+1+4\kappa)^{1-a}}{T-t}+\frac{(t-\tau+1+4\kappa)^{1-a}}{T-\tau+2+8\kappa}$
$\lesssim\frac{1}{\mu^{a}(T+\kappa)}\left((T+\kappa)^{1-a}(1+\log T)+(T+\kappa)^{1-a}\right)\lesssim\frac{1+\log T}{\mu^{a}(T+\kappa)^{a}}.$ (46)
Plug (45) and (46) back into (44) to have
$\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{a}}{\sum_{s=t}^{T-1}\gamma_{s}}\lesssim\frac{1+\log T}{\mu^{a}(T+\kappa)^{a}}.$
Applying the above inequality (for $a=1$ and $a=\mathsf{p}-1$) to (43), we conclude
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{HD^{2}}{\exp\left(\frac{T}{2+4\kappa}\right)}+\frac{G^{2}\left(1+\log T\right)}{\mu\left(T+\kappa\right)}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\left(1+\log T\right)}{\mu^{\mathsf{p}-1}\left(T+\kappa\right)^{\mathsf{p}-1}}.$
@@ E.6.2 Useful Existing Results
Note that Lemma 4 can be only applied to a randomized algorithm with bounded outputs. To overcome this issue, we need the following variants of $\bar{F}_{T,U}$ and $\bar{\boldsymbol{g}}_{T,U}$ introduced by [7]:
@@
    Let $\boldsymbol{u}_{j_{t}}\in\mathbb{R}^{d}$ denotes the $j_{t}$-th column of $U$. By the definition of $j_{t}$, there is
@@ E.6.3 Analysis under Heavy-Tailed Noise
The instances $\widehat{F}_{T,U}$ and $\widehat{\boldsymbol{g}}_{T,U}$ (see (59) and (60)) constructed based on $F_{T}$ in Lemma 7 and $\boldsymbol{g}_{T}$ in Lemma 8 under $\eta=\frac{1}{\sqrt{5}}-\frac{39}{140}$ and $R=7\gamma\sqrt{T}$ for $\gamma=23$ satisfy the following properties:
@@ F.1 Full Results for Smooth $\ell_{t}$
When $\boldsymbol{x}^{\star}\in\mathcal{X}$ is also a global minimizer (i.e., $\boldsymbol{x}^{\star}\in\mathrm{arginf}_{\boldsymbol{x}\in\mathbb{R}^{d}}F(\boldsymbol{x})$), we have $\left\|\nabla F(\boldsymbol{x}^{\star})\right\|=0$ since $F$ is differentiable. This fact recovers Corollary 4.
@@ F.2.2 New Regret for $\mathsf{OAdaGrad}$
Equipped with Lemma 10, we first prove the following Theorem 14, which establishes the regret of $\mathsf{OAdaGrad}$ under heavy tails.
@@
To gain some intuition of the new Condition 2, we first let $H_{t}=0$ to have $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim G_{t}$, which is similar to $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{y})\right\|\lesssim G_{t},\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$ in Assumption 3. Next, we consider the case $G_{t}=0$, where Condition 2 states an inequality $\left\|\nabla\ell_{t}(\boldsymbol{x})-\nabla\ell_{t}(\boldsymbol{x}_{t}^{\star})\right\|\lesssim H_{t}^{\frac{1}{1+\nu}}(\ell_{t}(\boldsymbol{x})-\ell_{t}^{\star})^{\frac{\nu}{1+\nu}}$, which is known to hold if $\ell_{t}$ is $(H_{t},\nu)$-Hölder smooth on $\mathbb{R}^{d}$ and convex, which can be viewed as a generalization of Assumption 3 (though strictly speaking, we only require $\ell_{t}$ to be Hölder smooth on $\mathcal{X}$ in Assumption 3). For general $G_{t}$, $H_{t}$, and $\nu$, whether Condition 2 is strictly general than the third point in Assumption 3 is unclear. Thus, we consider it as a separate condition here.
@@ G.2.2 Convergence Rates
Similar to before, we can consider two situations where problem-dependent parameters are known or unknown, corresponding to the following Corollaries 10 and 11, respectively. To save space, the proofs are omitted since they can be easily checked.
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim$ $\frac{\nabla}{N}+\frac{H\delta^{\nu}}{N^{\frac{1}{2}+\nu}}+\frac{G}{\sqrt{N}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\Delta}{\delta N}$
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim$ $\frac{\nabla}{N}+\frac{H\delta^{\nu}}{N^{\frac{1}{2}+\nu}}+\frac{H}{\delta^{\frac{2-\nu}{5}}N^{\frac{1+2\nu}{5}}}$
@@ G.2.3 A Special Case: Hölder Smooth Nonconvex Functions
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{KT}+\frac{\Delta}{\delta K}+\frac{\sigma}{T^{1-\frac{1}{\mathsf{p}}}}+H\delta^{\nu}.$
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{N}+\frac{H^{\frac{1}{1+\nu}}\Delta^{\frac{\nu}{1+\nu}}}{N^{\frac{\nu}{1+\nu}}}+\frac{\sigma}{N^{1-\frac{1}{\mathsf{p}}}}+\frac{\sigma^{\frac{\mathsf{p}\nu}{r}}H^{\frac{\mathsf{p}-1}{r}}\Delta^{\frac{(\mathsf{p}-1)\nu}{r}}}{N^{\frac{(\mathsf{p}-1)\nu}{r}}}.$
@@
$\mathbb{\mathbb{E}}\left[\frac{1}{K}\sum_{k=1}^{K}\left\|\nabla F(\bar{\boldsymbol{z}}_{k})\right\|_{\delta}\right]\lesssim\frac{\nabla}{N}+\frac{\Delta}{N^{\frac{1}{4}}}+\frac{H}{N^{\frac{\nu}{4}}}+\frac{\sigma}{N^{\frac{\mathsf{p}-1}{2\mathsf{p}}}}.$
@@
For general $\nu\in\left(0,1\right)$, as far as we know, no previous works consider heavy-tailed noise. Hence, both corollaries are the first and new.